6xsq.(2x+3)-5x(2x+3)=4(2x+3)
Answer by mananth(16946)

6x^2*(2x+3)-5(2x+3)=4(2x+3)
let 2x+3=a
6ax^2-5ax=4ax
6ax^2-5ax-4a=0
a(6x^2-5x-4)=0
either a=0
or (6x^2-5x-4)=0
If a= 0
2x+3=0
x=-3/2
if 6x^2-5x-4=0
6x^2-8x+3x-4=0
2x(3x-4)+1(3x-4)=0
(3x-4)(2x+1)=0
3x-4 =0
Or
2x+1=0
x=4/3
OR
x=-1/2
x=-1/2, or -3/2, or 4/3